OpenStudy (anonymous):

find an equation of the tangent line to teh curve y=(xsin(x))/(1+x) at the point (0,0).

OpenStudy (dumbcow):

tangent line: \[y - y_1 = \frac{dy}{dx}(x-x_1)\] (x1,y1) = (0,0) so \[y = \frac{dy}{dx} x\]

OpenStudy (dumbcow):

to find derivative use product and quotient rules

OpenStudy (dumbcow):

f = x*sin(x) f' = sin(x) + xcos(x) g = 1+x g' = 1 \[(\frac{f}{g})' = \frac{f'g - fg'}{g^{2}}\]

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.